The Soma Inn is trying to determine its break-even point. The inn has 75 rooms that are rented at $56 a night. Operating costs are as follows.

Salaries $10,600 per month
Utilities 2,600 per month
Depreciation 1,400 per month
Maintenance 700 per month
Maid service 7 per room
Other costs 31 per room

Correct answer iconYour answer is correct.

Determine the inn’s break-even point in (1) number of rented rooms per month and (2) dollars.

1. Break-even point in rooms
2. Break-even point

$

eTextbook and Media

Partially correct answer iconYour answer is partially correct.

If the inn plans on renting an average of 50 rooms per day (assuming a 30-day month), what is (1) the monthly margin of safety in dollars and (2) the margin of safety ratio? (Round ratio to 0 decimal places, e.g. 10.)

1. Margin of safety

$

2. Margin of safety ratio %
Answer #1

1) Break even point = 15300/(56-38) = 850 rooms

Break even point = 850*56 = $47600

2) Margin of safety in dollars = (1500-850)*56 = 36400

Margin of safety ratio = 36400/84000 = 43%